OpenStudy (anonymous):

How do I solve √8 - 15√12 ?

OpenStudy (anonymous):

√8 - 15√12 is the same as \[\sqrt{2^{2}*2} - 15\sqrt{2^2*3}\]

OpenStudy (anonymous):

then bring out the squares \[2\sqrt{2}-30\sqrt{3}\]

OpenStudy (anonymous):

>__< I'm sorry I meant " √8 - 15√2 ?"

OpenStudy (anonymous):

okay well that is easier \[\sqrt{2^{2}*2} - 15\sqrt{2}\] reduces to \[2\sqrt{2} - 15\sqrt{2}\] now subtract the coefficients

OpenStudy (anonymous):

Aah. So then it'd be a negative right. -13√2?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Ah. Thanks. :)